# Some configuration results for area-minimizing cones

### v4

@@ 4. Configuration results
In this section we will give a proof of Theorem 1.1 first which applies to arbitrary closed minimal submanifolds in Euclidean spheres. A further glance actually leads us to more results, such as Corollary 4.2 and Theorem 1.2. After that, we will move to some configuration results within the category of Type-c cones.
4.1. For general closed minimal submanifold.
@@
Therefore, when $n$ is large enough, by comparing leading term of (4.1) and (4.2) we have
@@
Note that, by [L-Z], for fixed $k_{1}\geq 3$ and every $\delta,\delta^{\prime}>0$ there exists spiral minimal product type $L_{1}$ of dimension $k_{1}$ in some sphere with its normal radius $R_{1}<\delta$ and $\frac{\alpha_{1}^{2}}{k_{1}}>(\delta^{\prime})^{-1}$ where $\alpha_{1}$ is its supremum norm of second fundamental form with respect to unit normals. So generally there is no control no these two quantities for a minimal submanifold (of codimension larger than two) in sphere. However, machinery Theorem 1.1 works regardless of the initial figure of $L_{1}$ at all.
Apparently, the idea here can lead to many general configuration results. Among others, here we would like to mention a corollary of the above proof.
Corollary 4.2.
@@
There exists integer $d_{0}$ such that the cone over the minimal product $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{\ell}}$ is area-minimizing whenever $d=\sum_{i=1}^{\ell}d_{i}\geq d_{0}$. If $L$ is a minimal isoparametric hypersurface or a focal submanifold, then this essentially falls into Theorem 4.18 in below and $k+d\geq 36$ would be sufficient for the area-minimality.
Remark 4.4.
Given a closed minimal hypersurface $L$ in $\mathbb{S}^{n}$ ($n\geq 3$) and any integer $\tau\geq 2$. By the corollary and remark, it follows that based on $L$ there exist infinitely many area-minimizing cones in Euclidean spaces having fixed codimension $\tau$.
Now we are ready to give a proof of Theorem 1.2.
Proof of Theorem 1.2. Set
@@
is injective.
Conversely, for every $(\subset\mathbb{S}^{N_{A}}) 𝒜\in\mathscr{A}$, according to Corollary 4.2 there exists a smallest integer $d_{A}\geq\dim A+1$ such that the cone over $A\times\mathbb{S}^{d}$ is area-minimizing when $d\geq d_{A}$. So, if we define
@@
Here we add $N_{A}$ in the dimensional part of $\mathbb{S}$ in (4.6) to distinguish $A$ being a minimal submanifold in different ambient spheres, since we regard $A\subset\mathbb{S}^{N_{A}}$ and $A\subset\mathbb{S}^{N_{A}+1}$ as two distinct elements in $\mathscr{A}$.
@@
In order to get more configuration results in the category of Lawlor cones, we dive into more details about the minimal product construction and the vanishing angle controls in the rest part of the paper.
@@
We wish to show that, in some circumstances, the minimal product of links $L_{1},L_{2}$ of two Type-c cones can again span a Type-c cone. So we aim to get the inequality $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ where $k=k_{1}+k_{2}$ and $k_{i}$ is the dimension of $L_{i}$ for $i=1,2$.
In this subsection we give some structural inequality for normal radius of minimal products. The first three results actually work for constant products, see Remark 3.3. Here the constant product construction merely means that $\{\lambda_{i}\}$ in the algorithm (3.1) are nonzero numbers satisfying $\sum_{i=1}^{n}\lambda_{i}^{2}=1$.
@@
since this would imply that $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ as we want. But, comparing this with Lemma 2.3, one can see that the ratio outside $\tan$ is $\frac{k_{1}+1}{k_{1}+k_{2}+1}$ which is larger than the constant $\frac{k_{1}}{k_{1}+k_{2}}$. It turns out that it could be very hard or likely impossible to improve the ratio on the right hand side of (2.4) in Lemma 2.3 to be $\frac{k_{1}}{k_{1}+k_{2}}$. (The same problem exists for the Type-F estimate (2.5) of Lemma 2.4.) To conquer this difficulty, instead we turn to improve the inequality (4.8) and get what we need for our purpose.
@@
As in Corollary 4.9, $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. Assume that (1) $k_{2}\geq 4$ or (2) $k_{2}\geq 3,k_{1}\geq 2$ and assume that $R_{1}\leq\frac{\pi}{2}$. Then for the minimal product construction we have
@@
$\leq$ $2k_{1}k_{2}(1-\cos R_{1})(1+\cos R_{1})+k_{1}^{2}\sin^{2}R_{1}$
@@
Now, following the argument of Corollary 4.9, in order for
@@
4.3. Configuration results in the category of Typc-c Lawlor cones.
For better exposition we introduce the following concepts.
@@
There are plentiful examples for each type. For some neat ones, we may consider isoparametric foliations of spheres. For brief information about their s(L)(L), readers are referred to [TZ]. Interested readers may check references therein for further knowledge.
@@
Example 4. Our Theorem 1.1 can manufacture a plethora of elements of class (b) and class (c). For example, given finitely many closed minimal submanifolds $\{L_{i}\}$ in spheres, whenever they are not much curved, i.e., $\alpha_{i}^{2}\leq k_{i}$, our Theorem 1.1 can apply and produce lots of members in class (b).
@@
A useful lemma for further discussion is the following.
@@
Remark 4.20.
By no means that $11$ is optimal, we simply take the advantage of the vanishing angle table of [Law]. One can definitely improve the sum requirement. Since successive minimal products make dimension larger and larger, so any concrete lower bound on total dimension would be fine for configuration results in below.
@@
Note that by Lemma 2.3, when $\alpha^{2}=k=k_{1}+k_{2}$,
@@
exists and, when $\alpha^{2}=\frac{k\alpha_{\tau}^{2}}{k_{\tau}}$ either for $\tau=1$ or for $\tau=2$,
@@
and the latter exists as by assumption either $C(L_{\tau})$ is a Type-c cone.
@@
Our first configuration result in this subsection is the following.
@@
Remark 4.22.
If $(L_{1})=s$$(L_{2})<1$, the proof in below will bother (4.15) with additional assumption $k_{1}+k_{2}\geq 11$ or some smaller number, see Remark 4.20.
@@
Case 1. $R_{1}\leq\frac{\pi}{2}$. Now we apply the key control Lemma 4.10 under the assumption on dimensions ($k_{1},\,k_{2}\geq 3$ would be enough for (4.3), as condition (2) of Lemma 4.10 satisfied) to get
@@
Case 2. $\frac{\pi}{2}<R_{1}\leq\pi$. By the proof of Lemma 4.19 for $(L_{1})=s$$(L_{2})\geq 1$ the vanishing angle $\theta_{c}(k+1,\alpha)$ always exists. Note that (4.15) is not bothered so that we have no assumption on total dimension. Moreover, there is a free rough control $\theta_{c}(k+1,\alpha)\leq\tan\theta_{c}(k+1,\alpha)<\frac{1}{\alpha}$ since the function $(1-\alpha t)e^{\alpha t}$ vanishes at $t=\frac{1}{\alpha}$ (see appendix for more information about the equivalent calibration inequality where solution needs to reach zero before $t=\frac{1}{\alpha}$).
@@
From the proof of Theorem 4.21, we know that as long as $\cos R$ and $\alpha^{2}$ are determined by the same slot of the minimal product then the minimal product spans a Type-c cone. There are many specific cases for which this coincidence in the same slot occurs. A general structure result is the following (cf. Corollary 4.2).
@@
Under the dimension assumption, we can see from (3.12) that
@@
Moreover, a further structure result related to Theorem 1.1 can be obtained.
@@
Suppose $L_{1}^{\times s_{1}}\times\cdots\times L_{n}^{\times s_{n}}$ with $n\geq 2$ and $s_{i}\geq 1$ for $i=1,\cdots,n$ spans a Type-c cone. Then so does $\tilde{L}\triangleq L_{1}^{\times(s_{1}+\ell_{1})}\times\cdots\times L_{n}^{\times(s_{n}+\ell_{n})}$ where $\ell_{i}\geq 0$ and $\sum_{i+1}^{n}k_{i}\ell_{i}\geq 3$.
@@
Here $L_{i}$ itself may not be area-minimizing.
@@
This result in particular indicates that, for given $\{L_{i}\}_{i=1}^{n}$, Theorem 1.1 can guarantee the existence of uniform $\tilde{k}$ such that whenever total dimension is no less than $\tilde{k}$ the cone C(L~)C($\tilde${L}) is then a Type-c cone and thus area-minimizing. When each $k_{i}\geq 3$, this $\tilde{k}$ could be less than $k$ in Remark 4.1 in general.
@@
Without loss of generality, we consider $\check{L}\triangleq L_{1}^{\times(s_{1}+1)}\times\cdots\times L_{n}^{\times s_{n}}$ with $k_{1}\geq 3$ only and other situations can be derived accordingly. Note that by Lemma 4.7 it follows $R(L)\leq\frac{\pi}{2}$ where $L\triangleq L_{1}^{\times s_{1}}\times\cdots\times L_{n}^{\times s_{n}}$. By Proposition 4.14 or Lemma 3.1 we have ${\tt s}(L)\geq\max\{1,{\tt s}(L_{1})\}$. Therefore, both $R(\check{L})$ and ${\tt s}(\check{L})$ are decided by the first input $L$ of the minimal product of $L$ and $L_{1}$. When $k_{1}\geq 3$, the key control Lemma 4.10 applies and the argument in the proof in Case 1 of Theorem 4.21 leads us to the conclusion of the theorem. ∎
In fact more subtle structures can be established. To illustrate the point, let us mention merely two more situations, based on which interested readers may figure out many more structures.
@@
Let $L^{k}$ be the link of a Type-c cone with $k\geq 3$. If $L$ is of class (II) with no other assumption on dimension or of class (a) with $d=k+\sum_{i=1}^{n}d_{i}\geq 11$, then $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{n}}$ spans a Type-c cone.
@@
The reason is following. If some slot $\mathbb{S}^{d_{\tau}}$ decides $R$ instead of component $L$, then we are in Case 2 in the proof of Theorem 4.21 and can get $\frac{R}{2}\geq\sqrt{\frac{d_{\tau}}{d}}$ now and $\theta_{c}(d+1,\alpha)<\frac{1}{\alpha}\leq\sqrt{\frac{d_{\tau}}{d}}$ by the existence of $\theta_{c}$ when $d\geq 11$. If $L$ is of class (II), the requirement on total dimension can be dropped. ∎
Instead of successively multiplying spheres, one can also consider using other kinds of links of class (I) (or their mixed types) for universal configuration results. For example, let $\tilde{M}$ be a focal submanifolds with ${\tt g}=4$ of sufficiently large dimension $n$. Since normal radius of each of them is always $\frac{\pi}{2}$, we get the following.
@@
One can take $L$ to be a resulting Type (II) cone from Theorem 1.1. Then Theorems 4.25, 4.29 and 4.30 explore some extra structures. Moreover, if $L$ is of class (b), then so are the resultings in these theorems.
@@
We would like to point out that Lemma 4.19, Theorem 4.21, Remarks 4.23, Corollary 4.24, Theorems 4.25, 4.29 and 4.30 will also be true if “Type-c” is replaced by “Type-F”. The key point is to verify (4.16) and (4.18) with the extra factor from (2.5). We provide details in appendix.
@@
When “Type-c” is replaced by “Type-F”, the requirement the total dimension $\geq 11$ in Lemma 4.19, Remark 4.22 and Theorem 4.29 can be replaced by the total dimension $\geq 7$, since corresponding upper bound of replaced (4.15) exists as $\theta_{F}(8,\sqrt{7})<16^{\circ}$ according to Lawlor’s table in [Law]. Also see $(\ast)$ in appendix for remaining the requirement $k_{1},k_{2}\geq 3$” unchanged in the Type-F version of Theorem 4.21.
Configuration results of this paper can provide a great number of local models in [z0] for observing duality obstruction of calibrations in smooth category.

### v5

@@ 4. Configuration results
This section consists of three parts.
In §4.1 the general configuration results Theorems 1.1 and 1.3 and Corollary 1.5 will be proved. The main idea is to gain $\theta_{c}<\frac{R}{2}$ in the asymptotic sense based on Lemmas 2.3 and 3.2.
In §4.2 some useful controls especially a key inequality on normal radius of minimal product $L_{1}\times L_{2}$ will be given in Lemma 4.9. Such an inequality together with Lemma 2.3 can lead to $\theta_{c}<\frac{R}{2}$ if both $\theta_{c}$ and $R$ of $L_{1}\times L_{2}$ are decided by the same component, which itself spans a Type-c cone. As a result, $L_{1}\times L_{2}$ also spans a Type-c cone.
In §4.3 we shall provide a partial positive answer to Question $(\star)$ in §1 under some assumptions. Besides, a quick proof of the main result of [TZ] will be explained in Theorem 4.17 following Remark 4.1. Moreover, we derive some more subtle structures (e.g. dimension gap of Theorem 1.1 in Theorem 4.23) complementing the asymptotic results in Theorems 1.1 and 1.3. Some other configuration results within the category of Type-c cones will also be discussed.
4.1. Asymptotic results for general closed minimal submanifold.
@@
Therefore, when $n$ is large enough, by comparing leading terms of (4.1) and (4.1) we have
@@
Note that, by [L-Z], for fixed $k_{1}\geq 3$ and every $\delta,\delta^{\prime}>0$ there exists spiral minimal product type $L_{1}$ of dimension $k_{1}$ in some sphere with its normal radius $R_{1}<\delta$ and $\frac{\alpha_{1}^{2}}{k_{1}}>(\delta^{\prime})^{-1}$ where $\alpha_{1}$ is its supremum norm of second fundamental form with respect to unit normals. So generally there is no control no these two quantities for a minimal submanifold (of codimension larger than two) in sphere. However, machinery works regardless of the initial figure of $L_{1}$ at all for spanning area-minimizing cones $C(L_{1}^{\times n})$ in the asymptotic sense in Theorem 1.1.
Apparently, the idea here can lead to many general configuration results. Among others, we get a proof of Theorem 1.3.
Proof of Theorem 1.3. Let $k_{1}$ be the dimension of $L\subset\mathbb{S}^{N}$. Then by Lemma 3.2 the normal radius $R=R(L\times\mathbb{S}^{d})$ of $L\times\mathbb{S}^{d}$ satisfies
$\cos R=1-\min\left\{\frac{k_{1}}{k_{1}+d}(1-\cos R_{1}),\,\frac{2d}{k_{1}+d}\right\}.$
Hence, when $d$ is large (e.g. $d\geq k_{1}$), we get
$\cos R=1-\frac{k_{1}}{k_{1}+d}(1-\cos R_{1})$
Similar to the argument in the proof of Theorem 1.1 we have
$\arccos\left(1-\frac{k_{1}}{k_{1}+d}(1-\cos R_{1})\right)=\frac{\sqrt{2k_{1}(1-\cos R_{1})}}{\sqrt{k_{1}+d}}+\text{ higher order terms}.$
Now note that by Lemma 3.1
$\alpha^{2}=(k_{1}+d)\max\left\{\frac{\alpha_{1}^{2}}{k_{1}},\,1\right\}.$
Therefore, according to Lemma 2.3 with $m=k_{1}+d+1$ and $r=12$, we get
$\tan\theta_{c}\Big(k_{1}+d+1,\alpha\Big)<\frac{12}{k_{1}+d+1}\tan\theta_{c}\left(12,\,\frac{12}{k_{1}+d+1}\sqrt{(k_{1}+d)\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}\right)$
when $k_{1}+d+1>12$.
So, when $d$ is large enough, by comparing leading terms we have
$\theta_{c}\left(k_{1}+d+1,\alpha\right)<\tan\theta_{c}\big(k_{1}+d+1,\alpha\big)<\frac{R(L\times\mathbb{S}^{d})}{2}.$
By Lawlor’s criterion Theorem 2.2 the cone $C(L\times\mathbb{S}^{d})$ is area-minimizing in $\mathbb{R}^{n+d+2}$.
If one replaces $\mathbb{S}^{d}$ by $\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{\ell}}$, then the generalization in Remark 1.4 holds due to the same kind of reasoning. $\square$
@@
Given a closed minimal hypersurface $L$ in $\mathbb{S}^{n}$ ($n\geq 3$) and any integer $\tau\geq 2$. By Theorem 1.3 and Remark 1.4, it follows that based on $L$ there exist infinitely many area-minimizing cones in Euclidean spaces having the fixed codimension $\tau$.
Beside the advantage of control on the precise codimension as in Remark 1.4, Theorem 1.3 can lead to a simple proof of Corollary 1.5.
Proof of Corollary 1.5. Set
@@
is obviously injective.
Conversely, for every $(\subset\mathbb{S}^{N_{A}}) 𝒜\in\mathscr{A}$, according to Theorem 1.3 there exists a smallest integer $d_{A}\geq\dim A+1$ such that the cone over $A\times\mathbb{S}^{d}$ is area-minimizing when $d\geq d_{A}$  [^4] Also see Theorem 4.23 for a possible gap behavior.. So, if we define
@@
Here we add $N_{A}$ in the dimensional part of $\mathbb{S}$ in (4.6) to distinguish $A$ being a minimal submanifold in different ambient spheres, since we regard $A\subset\mathbb{S}^{N_{A}}$ and $A(\subset\mathbb{S}^{N_{A}})\subset\mathbb{S}^{N_{A}+1}$ as two distinct elements in $\mathscr{A}$.
@@
Remark 4.5.
In fact one can also apply Theorem 1.1 for Corollary 1.5 but needs to be more careful in choosing corresponding $\tt G$ map.
Before closing this subsection, we would like to remark that all previously discovered area-minimizing cones C(L~)C($\tilde${L}) through Lawlor’s curvature criterion share the property (or restriction) that the curvature function $\alpha$ remains constant in the entire $\tilde{L}$, whereas in our Theorems 1.1 and 1.3 $\alpha$ can be a non-constant function over $\tilde{L}$.
@@
Instead of asymptotic behaviors in §4.1, we shall focus on some control regarding Question $(\star)$ of §1 in this subsection. More precisely, we wish to show that, in some circumstances, the minimal product of links $L_{1},L_{2}$ of two Type-c cones can again span a Type-c cone. So we aim to get the inequality $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ where $k=k_{1}+k_{2}$ and $k_{i}$ is the dimension of $L_{i}$ for $i=1,2$.
In particular we will give some structural inequality for normal radius of minimal products. The first three results actually work for constant products, see Remark 3.3. Here the constant product construction merely means that $\{\lambda_{i}\}$ in the algorithm (3.1) are nonzero numbers satisfying $\sum_{i=1}^{n}\lambda_{i}^{2}=1$.
@@
since this could imply that $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ as we want. But, comparing this with Lemma 2.3, one can see that the ratio outside $\tan$ is $\frac{k_{1}+1}{k_{1}+k_{2}+1}$ which is larger than the constant $\frac{k_{1}}{k_{1}+k_{2}}$. It seems very hard or likely impossible to improve the ratio on the right hand side of (2.4) in Lemma 2.3 to be $\frac{k_{1}}{k_{1}+k_{2}}$. (The same problem exists for the Type-F estimate (2.5) of Lemma 2.4.) To conquer this difficulty, instead we turn to improve the inequality (4.8) and get what we need for our purpose.
@@
As in Corollary 4.8, $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. Assume
(1) $k_{2}\geq 4$ or (2) $k_{2}\geq 3,k_{1}\geq 2$
and assume $R_{1}\leq\frac{\pi}{2}$. Then for the minimal product construction we have
@@
$\left(\text{by assumption }0<R_{1}\leq\frac{\pi}{2}\right)\ \ \ \ \ \ \displaystyle\leq$ $2k_{1}k_{2}(1-\cos R_{1})(1+\cos R_{1})+k_{1}^{2}\sin^{2}R_{1}$
@@
Following the argument of Corollary 4.8 but now with $\sin R\leq\frac{\sqrt{k_{1}(2k_{2}+k_{1})}}{k}\sin R_{1}$, in order for
@@
With the preparation of this subsection, we are ready to derive more configuration results in the next subsection.
4.3. More configuration results.
For better exposition, let us introduce some useful concepts.
@@
There are plentiful examples for each type. For some neat ones, we may consider isoparametric foliations of spheres. Here we mention some brief information about their s(L)(L). Interested readers may check [TZ] and references therein for further knowledge.
@@
Example 4. Our Theorem 1.1 can manufacture a constellation of elements of class (b) and class (c). For example, given finitely many closed minimal submanifolds $\{L_{i}\}$ in spheres, whenever they are not much curved, i.e., $\alpha_{i}^{2}\leq k_{i}$, our Theorem 1.1 can apply and produce lots of members in class (b).
@@
Now let us prove a useful lemma on the existence of vanishing angle.
@@
When $\alpha^{2}=k=k_{1}+k_{2}$, by Lemma 2.3 one can see that
@@
exists.
When $\alpha^{2}=\frac{k\alpha_{\tau}^{2}}{k_{\tau}}$ either for $\tau=1$ or for $\tau=2$,
@@
and thus $\theta_{c}(k+1,\alpha)$ exists by the assumption $C(L_{\tau})$ is a Type-c cone.
@@
We are ready to obtain more configuration results in this subsection.
@@
Case 1. $R_{1}\leq\frac{\pi}{2}$. In this case we can apply the key control Lemma 4.9 under the assumption on dimensions ($k_{1},\,k_{2}\geq 3$ would be enough for (4.3), as condition (2) of Lemma 4.9 satisfied) to get
@@
Case 2. $\frac{\pi}{2}<R_{1}\leq\pi$. By the proof of Lemma 4.18 for $(L_{1})=s$$(L_{2})\geq 1$ the vanishing angle $\theta_{c}(k+1,\alpha)$ always exists without bothering (4.15) so that we have no assumption on total dimension. Moreover, there is a free rough control $\theta_{c}(k+1,\alpha)\leq\tan\theta_{c}(k+1,\alpha)<\frac{1}{\alpha}$ since the function $(1-\alpha t)e^{\alpha t}$ vanishes at $t=\frac{1}{\alpha}$ (see appendix for more information about the equivalent calibration inequality where solution needs to reach zero before $t=\frac{1}{\alpha}$).
@@
From the proof of Theorem 4.19, we know that as long as $\cos R$ and $\alpha^{2}$ are determined by the same slot of the minimal product then the minimal product spans a Type-c cone. There are many specific cases for which this coincidence in the same slot occurs. A general structure result is the following (cf. Theorem 1.3).
@@
Under the dimension assumption, we can see from Lemma 3.2 that
@@
Moreover, a subtle structure result related to Theorem 1.1 can be obtained.
@@
Suppose $L=L_{1}^{\times s_{1}}\times\cdots\times L_{n}^{\times s_{n}}$ with $n\geq 2$, $s_{i}\geq 1$ and each component dimension $k_{i}>0$ for $i=1,\cdots,n$ spans a Type-c cone. Then so does $\tilde{L}\triangleq L_{1}^{\times(s_{1}+\ell_{1})}\times\cdots\times L_{n}^{\times(s_{n}+\ell_{n})}$ where $\ell_{i}\geq 0$ and $\sum_{i+1}^{n}k_{i}\ell_{i}\geq 3$.
@@
Here $L_{i}$ itself may not span area-minimizing cone.
@@
If we use $L\prec\check{L}$ to describe the relation of increasing exponential order between $L$ and $\check{L}$, then, given a strictly increasing sequence of this relation, according to Remark 4.1 there must be a (largest) successive infinite tail elements of which all span area-minimizing cones. The result here indicates that the tail and the element in the sequence which spans an area-minimizing cone of lowest dimension has gap at most $2$ in dimension.
@@
Without loss of generality, we consider $\check{L}\triangleq L_{1}^{\times(s_{1}+1)}\times\cdots\times L_{n}^{\times s_{n}}$ with $k_{1}\geq 3$ only and other situations can be derived accordingly. Note that by Lemma 4.6 it follows $R(L)\leq\frac{\pi}{2}$. By Proposition 4.13 or Lemma 3.1 we have ${\tt s}(L)\geq\max\{1,{\tt s}(L_{1})\}$. Therefore, both $R(\check{L})$ and ${\tt s}(\check{L})$ are decided by the first input $L$ of the minimal product of $L$ and $L_{1}$. When $k_{1}\geq 3$, the key control Lemma 4.9 applies and the argument in the proof in Case 1 of Theorem 4.19 leads us to the conclusion of the theorem. ∎
In fact more subtle structures can be established. To illustrate the point, let us mention two more situations, based on which interested readers may figure out many others.
@@
Let $L^{k}$ be the link of a Type-c cone with $k\geq 3$. If $L$ is of class (II) then $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{n}}$ spans a Type-c cone.
@@
The reason is following. If some slot $\mathbb{S}^{d_{\tau}}$ decides $R$ instead of component $L$, then we are in Case 2 in the proof of Theorem 4.19 and can get $\frac{R}{2}\geq\sqrt{\frac{d_{\tau}}{d}}$ now and $\theta_{c}(d+1,\alpha)<\frac{1}{\alpha}\leq\sqrt{\frac{1}{d}}\leq\sqrt{\frac{d_{\tau}}{d}}$ by the existence of $\theta_{c}$.
When $R$ is decided by $L$, one only needs to note that $\alpha$ is also decided by $L$ and the statement follows according to the proof of Theorem 4.19. ∎
Instead of successively multiplying spheres, one can also consider using other kinds of links of class (I) (or their mixed types). For example, let $\tilde{M}$ be a focal submanifolds with ${\tt g}=4$ of sufficiently large dimension $n$. Since normal radius of each of them is always $\frac{\pi}{2}$, we get the following.
@@
Proof.
By the same kind of reasoning in the proof of Theorem 4.22. ∎
@@
One can take $L$ to be a resulting Type (II) cone from Theorem 1.1. Then Theorems 4.22, 4.26 and 4.27 explore some extra structures. Moreover, if $L$ is of class (b), then so are the resultings in these theorems.
@@
We would like to point out that Lemma 4.18, Theorem 4.19, Remarks 4.20, Corollary 4.21, Theorems 4.22, 4.26 and 4.27 will also be true if “Type-c” is replaced by “Type-F”. The key point is to verify (4.16) and (4.18) with the extra factor from (2.5). We provide details in appendix.
@@
When “Type-c” is replaced by “Type-F”, the requirement the total dimension $\geq 11$ in Lemma 4.18 and Theorem 4.26 can be replaced by the total dimension $\geq 7$, since corresponding upper bound of replaced (4.15) exists as $\theta_{F}(8,\sqrt{7})<16^{\circ}$ according to Lawlor’s table in [Law]. Also see $(\ast)$ in appendix for remaining the requirement $k_{1},k_{2}\geq 3$” unchanged in the Type-F version of Theorem 4.19.
Based on the configuration results of this paper, in [z0] we get a constellation of area-minimizing cones of codimension larger than one which cannot be calibrated by any globally defined continuous calibrations.